# Frostman random variables, entropy inequalities, and applications

### v4

@@ 4.4 More general polynomials
We begin by recording the results, or slight variants thereof, that we use from Shmerkin and Wang, [24], as well as the requisite material to state them. We then use these to explain how to prove Theorem 5.2. Our proof basically runs parallel to the proof of their Theorem 5.2, but with slightly different hypotheses and a different conclusion.
@@ 5.2.1 Material from Shmerkin and Wang
While it is clear that if $(\mu,\nu)$ have $(t,K,c,\delta)$-strong thing tubes, they have $(t,K,c,\delta)$-thin tubes, the following result from [24] shows that a partial converse also holds.
Lemma 5.5.
[Lemma 5.9 in [24]] Suppose $(\mu,\nu)$ have $(t,K,1-c,\delta)$-thin tubes. Then they have $(t,K^{\prime},c^{\prime},\delta)$-strong thin tubes, where $c^{\prime}>1-3c$, $K^{\prime}=2K$ if $c<1/4$, and $c^{\prime}>(1-\sqrt{c})/2$, $K^{\prime}=(1-\sqrt{c})^{-1}K$ if $c\geq 1/4$.
In [24], they use a result due to Orponen [20], which they call Proposition 5.11. We need something slightly different, so we state and prove explicitly what we will use here. For $v\in\mathbb{R}^{2},$ we will define $P_{v}$ as $P_{v}(x)=x\cdot v,$ so $P_{v}\mu$ is the pullback of $\mu$ through the map $P_{v},$ which means
@@
Proposition 5.6.
[Variant of Proposition 5.11 in [24]] Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in [0,1],[0,1],then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
@@
Lemma 5.7.
[Lemma 5.16 in [24]] Let $\rho\in\mathcal{P}(S^{1})$ be a $(\delta,s,K)$-measure and fix $\epsilon>0$. If $\delta<K^{-C^{\prime}\epsilon^{-1}}$ for some absolute constant $C^{\prime}$, then $\eta(s,t)$ can be chosen so that $\rho$ is $(D_{s}-\epsilon;1)$-adapted at scale $\delta$.
@@
Lemma 5.8.
[Lemma 4.7 from [24]] Fix $\tau$ and $D$ as above. If $D$ is (Lipschitz) continuous, then $s\mapsto\Sigma_{\tau}(D;s,d)$ is (locally Lipschitz) continuous on $(0,d)$.
@@
Theorem 5.9.
[Variant of Theorem 4.8 in [24]] Let $1\leq k\leq d,$ be natural numbers, $D:[0,d]\to[0,k]$ be a continuous function. Fix $c\in(0,1)$, $C\geq 1$, $0<\delta_{1}\leq 1$. Let $(Y,\nu)$ be a compact measure space, $U\subseteq\mathbb{R}^{d}$ be a bounded domain, and let $F_{y}:U\rightarrow\mathbb{R}^{k}$ be either the radial projection to $y$ or the distance to $y$.
@@
This is a technical result taken directly from Shmerkin and Wang.
Lemma 5.10.
Let $\rho\in\mathcal{P}(S^{1})$ be a $(\delta,s,K)$-measure and fix $\epsilon>0$. If $\delta<K^{-C^{\prime}\epsilon^{-1}}$ for some absolute constant $C^{\prime}$, then $\eta(s,t)$ can be chosen so that $\rho$ is $(D_{s}-\epsilon,1)$-adapted at scale $\delta.$
@@
Proposition 5.11.
[Proposition 5.17 in [24]] Given $\zeta>0$ and $u\in(0,1],$ there are $\xi>0$ and $\tau>0,$ both depending on $u$ and $\zeta,$ such that if $s\in(0,\varphi(u)-\zeta]$ then $\Sigma_{\tau}(D_{s};u,2)\geq s+\xi.$
@@
Lemma 5.12.
[Variant of Lemma 5.18 in [24]] Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has (s,K,1c)(s,K,1-c)-strong thin tube for some number $K>0,$ then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
@@
Next, we apply Lemma 5.4 to $\rho_{y}$ to obtain that for each $y$ there is a set $X^{\prime}_{y}\subseteq S_{1}$ for which $(\rho_{y})_{X^{\prime}_{y}}(B(x,r))<r^{s+\eta}$ for every $x<\delta_{0}$ and
@@
5.2.2 Modifications to their argument
@@
Finally, using (82), and Lemma 2.32, we can calculate the entropy as
@@
$\geq(1-c^{\prime})^{3}(n(\varphi(2s)-\zeta/2)-\log(\delta_{0})^{-1}).$
Since $\delta_{0}^{-1}=C_{1}\min\{K,\zeta^{-1}\}^{C_{1}}$ and $K=(C^{\prime}\min\{c_{0}^{-1},C\})^{C^{\prime O(1/\eta)}},$ we have that $\log\delta_{0}^{-1}=C_{2}\log C+C_{2}$ for some large constant $C_{2}$. Furthermore we can choose $M$ large enough in terms of $\zeta$ and $s$ so that $(1-c^{\prime})^{3}(\varphi(2s)-\zeta/2)>(1-2\zeta/M)^{3}(\varphi(2s)-\zeta/2)>(\varphi(2s)-\zeta),$ which concludes the proof. ∎
@@
Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$. Let $\epsilon=\epsilon(s)$ to be determined later.
Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$. We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$. Then
@@
When $\delta<r_{2}\leq 1$, the concentration conditions gives
@@
By Theorem 1.2’ (see Section 4.3), there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
@@
Putting (83), (84), and (86)- (88) together, one arrives at
@@
whenever $\delta$ is sufficiently small depending on $\phi$ and $s$ and $C$. The proof is completed. ∎
@@
Similar arguments as above will be applied. We only mention the differences that is needed to make.
@@
for some $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)>0$, which results from Theorem 1.3’ (see Section 4.4). Moreover, we substitute
@@
We use similar arguments as in the proof of Theorem 1.5, and work with $G=B\times B$. Since $B\geq\delta^{\epsilon}\#A\geq\delta^{-s+\epsilon}$, one has $\mathcal{L}^{1}(B_{\eta\delta})\gg\delta^{1-s+\epsilon}$. Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}=B_{\eta\delta}\times B_{\eta\delta}$. It is not hard to see that $X$ and $Y$ are i.i.d.. Combining the non-concentration condition, one deduces that
@@
As a replacement of (86), Theorem 1.4’ (see Section 4.4) shows that
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{\phi,s}(1)$
for some $\epsilon^{\prime}=\epsilon^{\prime}(s)>0$ and for all $n\in\mathbb{N}$. And the similar conclusion as in (90) appears that

### v5

@@ 4.4 More general polynomials
We prove Theorem 5.2 by adapting parts of the framework of Shmerkin and Wang [35] to an entropy setting. To keep the exposition self-contained, we first recall the definitions and auxiliary results from [35] that we use, together with a few minor variants stated in the form convenient for us. After this, we follow the main line of their argument, but we work under slightly different hypotheses and our conclusion is an entropy lower bound rather than a robustness statement.
To guide the reader, we indicate what is imported and what is new. We import from [35] the notions of robust measures, thin tubes, and adaptedness, and we use the quantitative estimates in [35] built on these concepts.
What we modify is the bookkeeping of hypotheses and constants. In particular, we track the dependence on the Frostman constant $C$ throughout, in a form uniform in the conditioning parameters that appear later. What is new here is the final conversion step, where we turn the robustness information delivered by the thin-tube argument into a dyadic entropy lower bound at a fixed scale, with an explicit loss of size $O(\log C)$.
@@ 5.2.1 Material from Shmerkin and Wang
While it is clear that if $(\mu,\nu)$ have $(t,K,c,\delta)$-strong thin tubes, they have $(t,K,c,\delta)$-thin tubes, the following result from [35] shows that a partial converse also holds.
Lemma 5.5 (Lemma 5.9 in [35]).
Suppose $(\mu,\nu)$ have $(t,K,1-c,\delta)$-thin tubes. Then they have $(t,K^{\prime},c^{\prime},\delta)$-strong thin tubes, where $c^{\prime}>1-3c$, $K^{\prime}=2K$ if $c<1/4$, and $c^{\prime}>(1-\sqrt{c})/2$, $K^{\prime}=(1-\sqrt{c})^{-1}K$ if $c\geq 1/4$.
In [35], they use a result due to Orponen [25], which they call Proposition 5.11. We need something slightly different, so we state and prove explicitly what we will use here. For $v\in\mathbb{R}^{2},$ we will define $P_{v}$ as $P_{v}(x)=x\cdot v,$ so $P_{v}\mu$ is the pullforward of $\mu$ through the map $P_{v},$ which means
@@
Proposition 5.6 (Variant of Proposition 5.11 in [35]).
Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in $[0,1],$ then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
@@
Lemma 5.7 (Lemma 5.16 in [35]).
Let $\rho\in\mathcal{P}(S^{1})$ be a $(\delta,s,K)$-measure and fix $\epsilon>0$. If $\delta<K^{-C^{\prime}\epsilon^{-1}}$ for some absolute constant $C^{\prime}$, then $\eta(s,t)$ can be chosen so that $\rho$ is $(D_{s}-\epsilon;1)$-adapted at scale $\delta$.
@@
Lemma 5.8 (Lemma 4.7 from [35]).
Fix $\tau$ and $D$ as above. If $D$ is (Lipschitz) continuous, then $s\mapsto\Sigma_{\tau}(D;s,d)$ is (locally Lipschitz) continuous on $(0,d)$.
@@
Theorem 5.9 (Variant of Theorem 4.8 in [35]).
Let $1\leq k\leq d$ be natural numbers, $D:[0,d]\to[0,k]$ be a continuous function. Fix $c\in(0,1)$, $C\geq 1$, $0<\delta_{1}\leq 1$. Let $(Y,\nu)$ be a compact measure space, $U\subseteq\mathbb{R}^{d}$ be a bounded domain, and let $F_{y}:U\rightarrow\mathbb{R}^{k}$ be either the radial projection to $y$ or the distance to $y$.
@@
Proposition 5.10 (Proposition 5.17 in [35]).
Given $\zeta>0$ and $u\in(0,1],$ there are $\xi>0$ and $\tau>0,$ both depending on $u$ and $\zeta,$ such that if $s\in(0,\varphi(u)-\zeta]$ then $\Sigma_{\tau}(D_{s};u,2)\geq s+\xi.$
@@
Lemma 5.11 (Variant of Lemma 5.18 in [35]).
Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has (s,K,1c)(s,K,1-c)-strong thin tube for some number $K>0$ and $c>0$, then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
@@
Next, we apply Lemma 5.4 to $\rho_{y}$ to obtain that for each $y$ there is a set $X^{\prime}_{y}\subseteq S^{1}$ for which $(\rho_{y})_{X^{\prime}_{y}}(B(x,r))<r^{s+\eta}$ for every $x<\delta_{0}$ and
@@
5.2.2 Modifications to their arguments
@@
Finally, since (99) hold for all $n$, and by Lemma 2.32, we can bounds the entropy as
@@
$\geq(1-c^{\prime})^{3}(n(\varphi(2s)-\zeta/2)-\log\delta_{0}^{-1}).$
Since $\delta_{0}^{-1}=C_{1}\min\{K,\zeta^{-1}\}^{C_{1}}$ and $K=(C^{\prime}\min\{c_{0}^{-1},C\})^{C^{\prime O(1/\eta)}},$ we have that $\log\delta_{0}^{-1}=C_{2}\log C+C_{2}$ for some large constant $C_{2}$. Furthermore we can choose $M$ large enough in terms of $\zeta$ and $s$ so that $(1-c^{\prime})^{3}(\varphi(2s)-\zeta/2)>(1-2\zeta/M)^{3}(\varphi(2s)-\zeta/2)>(\varphi(2s)-\zeta),$ which concludes the proof.
@@
Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$. Let $\epsilon=\epsilon(d,s)$ to be determined later.
Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipschitz on $0,1]^{2}$. We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$. Then
@@
When $\delta<r_{2}\leq 1$, the concentration condition gives
@@
By Theorem 1.2’ (see Section 4.3), there exists an $\epsilon^{\prime}=\epsilon^{\prime}(d,s)$ such that
@@
Putting (100), (101), and (103)–(105) together, one arrives at
@@
whenever $\delta$ is sufficiently small depending on $\phi$, $s$ and $C$. The proof is completed. ∎
@@
Arguments similar to those above will be applied. We only highlight the key differences.
@@
for some $\epsilon^{\prime}>0$, which results from Theorems 1.3’ (see Section 4.4). Moreover, we substitute
@@
This argument is also similar to that of the proof of Theorem 1.4, but here we work with $G=B\times B$. Since $B\geq\delta^{\epsilon}\#A\geq\delta^{-s+\epsilon}$, one has $\mathcal{L}^{1}(B_{\eta\delta})\gg\delta^{1-s+\epsilon}$. Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}=B_{\eta\delta}\times B_{\eta\delta}$. One checks that $X$ and $Y$ are i.i.d.. Combining the non-concentration condition, one deduces that
@@
As a replacement of (103), Lemma 4.8 and Theorem 1.3” (see Section 4.4) shows that
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)$
for some $\epsilon^{\prime}>0$ and for all $n\in\mathbb{N}$. Finally, a similar conclusion as in (107) appears, namely that